k(q)^2 / r = mv^2 please solve for v where k = 8.99x10^9 C, q = 10x10^-6 C, r = 0.055m, m = .001g. This problem is solving for electric potential energy and electric potential the right answer is V = 4.04x10^3 m/s. But I keep getting the wrong answer please do a step by step to show me thank you.
Answer #1

Solution)

Kq^2/r= mv^2

(8.99*10^9)*(10*10^-6)^2/0.055 =0.001*10^-3 *v^2

16.34= 10^-6 * V^2

So, V= 4042.27 = 4.04*10^3 m/s

========

Good luck:)